OpenStudy (anonymous):

how do you solve this?! 5+15/(n-3)=45/(n^2-3n) ???

OpenStudy (anonymous):

\[5+15/(n-3)=45/(n ^{2}-3n)\]

OpenStudy (anonymous):

n^2 - 3n = n(n-3)

OpenStudy (anonymous):

mutliply both side by (n-3) as well as by (n^2 - 3n).